OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of −12 ? y=x2−2 y=−x2−1 y=2x+1 y=−2x−1

OpenStudy (shamim):

required equation of a line is y=mx+c

OpenStudy (anonymous):

how would i put that in the equation?

OpenStudy (shamim):

given slope m =-12 x=-6 y=2

OpenStudy (shamim):

put those values in the above equation of a line find c

OpenStudy (radar):

There seems to be a conflict with the data you've provided for this problem. Review the problem and the distractors.

OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of −12 ? y=x2−2 y=−x2−1 y=2x+1 y=−2x−1 is the exact information in the problem

OpenStudy (radar):

The -12 seems to a stranger to this problem. If it is a -12 then the line equation would be: y = -12x - 70 which isn't even close to your options.

OpenStudy (radar):

You can substitute -6 for x and solve each option, the y value should compute as 2. y= -6^2 - 2 = 34 not 2 wrong answer for the first choice.

OpenStudy (anonymous):

so which one would that be... im not very good at math

OpenStudy (radar):

well lets check the 2nd one y= - (-6^2) - 1 = -36 -1 = -37 not the required 2. wrong answer1. Now the 3rd choice y = 2x + 1, y=2(-6) + 1 = -12 + 1 = 11, again not the 2 needed wrong answer. The last one y = -2x - 1 = -2(-6) -1 = 12 -1 = 11, again not the y =2 as provided by the point given for the problem (-6, 2) We can conclude there is no best answer for the choices provided.

OpenStudy (anonymous):

ok,that seems to be a problem

OpenStudy (radar):

The correct answer for the info provided is: y = -12x - 70 To verify substitute the point values (-6, 2) this means x = -6, and y = 2/ y=-12 (-6) -70 = 72 - 70 = 2, y=2 Yep correct answer.

OpenStudy (radar):

Did this problem come from your text, or a student handout?

OpenStudy (radar):

Sorry I can't be of further help. Good luck with your studies.

OpenStudy (radar):

By the way, the slope will be the coefficient of the "x" term in the equation provided by @shamim

OpenStudy (anonymous):

It came from a online school,you were a great help and thank you.

OpenStudy (radar):

You're welcome shadow.655